# Spatiotemporal Photon Blockade for Nonreciprocal Quantum Absorption

### v3

@@ Abstract
Controlling the flow of photons is crucial for advancing quantum technologies. We introduce the concept of spatiotemporal photon blockade for nonreciprocal quantum absorption, utilizing space-time-periodic metasurfaces. Our study presents a methodology for experimentally realizing this effect, where photon frequency coherence with the metasurface’s space-time modulation enables one-way quantum absorption. In this system, forward-traveling photons are energetically modulated and absorbed within the slab, while backward-traveling photons are transmitted without interaction. Our analysis includes band structure, isofrequency diagrams, and nonreciprocal absorption results. These findings lay the groundwork for developing nonreciprocal quantum devices and enhancing photon management in milli-Kelvin temperature quantum systems.
@@
Photon blockade is a quantum optical phenomenon analogous to the Coulomb blockade, where the suppression of electrical transport occurs due to electronic repulsion, preventing multiple electrons from occupying the same region. In photon blockade, the absorption or transmission of a single photon inhibits the absorption or transmission of subsequent photons. This phenomenon occurs in systems where strong nonlinear interactions between photons effectively create a scenario in which only one photon can be absorbed or transmitted at a time, leading to quantized light behavior. This effect, typically observed in systems such as atoms, quantum dots, or superconducting circuits coupled to resonant cavities, involves the system entering an excited state after the first photon interaction. This state shift prevents further photon interactions until the system returns to its ground state, which is crucial for quantum information processing and communication. Historically, photon blockade has been achieved in systems with strong coupling between quantum emitters and optical cavities, where the nonlinearity creates an energy gap sufficient to block additional photons  (?, ?). Recent advancements have expanded this concept to nonreciprocal photon blockade in rotating nonlinear devices, where device geometry or external fields control the directionality of photon interactions  (?). Additionally, loss-induced nonreciprocal photon blockade has been demonstrated, where engineered losses enable nonreciprocal photon dynamics, enhancing control over photon flow in quantum circuits  (?, ?).
@@
Quantum nonreciprocity is increasingly relevant for superconducting circuits, which are essential for scalable quantum computing and communication networks. Nonreciprocal elements like isolators and circulators are critical for protecting qubits from back-reflected signals and routing quantum information. Recent studies have explored various mechanisms for achieving quantum nonreciprocity, including dynamic modulation and parametric interactions in superconducting devices  (?, ?). Observations of superconducting nonreciprocity through diamagnetic effects and engineered circuit asymmetries offer robust platforms for nonreciprocal quantum devices  (?, ?, ?, ?). Traditional directional absorbers and electronic components, such as varactors, transistors, and diodes, face limitations and introduce noise at millikelvin temperatures typical of quantum technologies. To overcome these challenges, we propose a novel approach for nonreciprocal photon blockade through spatiotemporal modulation in quantum metasurfaces. By leveraging the unique properties of superconducting materials and dynamically modulating the metasurface structure, we achieve directional quantum absorption, where photons are selectively absorbed or transmitted based on their direction of incidence. This spatiotemporal photon blockade presents new opportunities for controlling quantum light-matter interactions and offers significant advancements for quantum information technologies and nonreciprocal photonic devices. Our approach achieves efficient directional absorption with simultaneous amplification, a rare combination in conventional devices. It addresses inefficiencies and spurious signal generation, paving the way for high-performance signal processing applications.
@@
Space-time refractive index modulation is conventionally realized through the modulation of electrical permittivity, as it is challenging to spatio-temporally modulate the magnetic permeability of materials  (?). In  (?, ?), it was shown for the first time that modulating both permeability and permittivity results in physical phenomena that cannot be achieved through conventional permittivity-modulated media, including zero reflection, zero band-gap, a large nonreciprocal response, and strong space-time harmonics. Furthermore, our metasurface demonstrates nonreciprocal behavior, with different absorption responses for incident waves from opposing directions, emphasizing the unidirectional properties imparted by the spatiotemporal modulation  (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?) and time-varying systems (?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?, ?). Utilizing analytical techniques such as Bloch-Floquet solutions, we provide a comprehensive understanding of the wave scattering phenomena from our metasurface, validating its design and offering insights for further optimization. Unlike linear systems constrained by phase-matching and dispersion, our nonlinear spatiotemporal media achieve superior performance in a compact form, making them ideal for quantum technologies.
1 Spatiotemporal Photon Blockade
Consider photons with frequency $\omega_{0}$ incident on a slab with a space-time periodic current density $J(z,t)$, modulated at a temporal frequency $\omega_{\text{s}}$, where $\omega_{\text{s}}=\omega_{0}$. We demonstrate that the temporal coherence between the incident photons and the periodic modulation of the superconducting slab induces a nonreciprocal photon blockade. This phenomenon arises because the energy associated with the space-time modulation effectively inhibits the transmission of subsequent photons. Instead, the incident photons transition to higher energy states within the slab and are absorbed, preventing their exit. This results in a one-way quantum absorption process, facilitated by the coherent spatiotemporal modulation of the superconducting slab.
@@ 1 Spatiotemporal Photon Blockade
Figure 1: Nonreciprocal photon blockade in superconducting spatiotemporal metasurfaces, where the space-time modulation and incident photons share the same frequency ($\omega_{\text{s}}=\omega_{0}$). The system comprises space-time-modulated nonlinear unit cells coupled together. a) Incidence from the left port ($L_{\text{in}}$): Photon blockade and energy transition to higher states occur due to the strong interaction and one-way coherency with the nonlinear left-to-right traveling space-time modulation, resulting in zero transmission to the right port ($R_{\text{out}}=0$). b) Illustration of the physical mechanism for incidence from the right port ($R_{\text{in}}$): Free passage of photons to the left port ($L_{\text{out}}=R_{\text{in}}$) occurs due to the lack of transition to higher energy states and weak interaction with the opposite-direction traveling nonlinear space-time modulation, resulting in full transmission of photons.

### v4

# Photon-Blockade Analogue Nonreciprocal Absorption in Spatiotemporal Metasurfaces

@@ Abstract
Controlling the flow of electromagnetic energy is essential for advancing quantum technologies. We introduce a spatiotemporally modulated superconducting metasurface that exhibits photon-blockade-analogue nonreciprocal absorption. In this system, the frequency of incident radiation is matched to the modulation frequency of the metasurface, enabling one-way directional absorption. Forward-traveling waves undergo resonant coupling to higher-order Floquet harmonics and are absorbed within the slab, while backward-traveling waves transmit freely without interaction. This behavior arises from classical wave interference and harmonic conversion in a space-time periodic medium—a classical analogue of quantum photon blockade. We present a design based on a superconductor-semiconductor metasurface incorporating cascaded Josephson field-effect transistors (JoFETs) for millikelvin-temperature operation. Our analysis includes the system Hamiltonian, Floquet band structure, isofrequency diagrams, and full-wave simulations demonstrating strong nonreciprocal absorption. These findings establish a pathway toward compact, nonreciprocal superconducting devices for quantum information processing and microwave photonics.
@@
Photon blockade is a quantum optical phenomenon where the absorption or transmission of a single photon inhibits the absorption or transmission of subsequent photons. This effect, typically observed in systems such as atoms, quantum dots, or superconducting circuits coupled to resonant cavities, involves the system entering an excited state after the first photon interaction. This state shift prevents further photon interactions until the system returns to its ground state, which is crucial for quantum information processing and communication. Historically, photon blockade has been achieved in systems with strong coupling between quantum emitters and optical cavities, where the nonlinearity creates an energy gap sufficient to block additional photons  (?, ?). Recent advancements have expanded this concept to nonreciprocal photon blockade in rotating nonlinear devices, where device geometry or external fields control the directionality of photon interactions  (?). Additionally, loss-induced nonreciprocal photon blockade has been demonstrated, where engineered losses enable nonreciprocal photon dynamics, enhancing control over photon flow in quantum circuits (?, ?).
@@
Quantum nonreciprocity is increasingly relevant for superconducting circuits, which are essential for scalable quantum computing and communication networks. Nonreciprocal elements like isolators and circulators are critical for protecting qubits from back-reflected signals and routing quantum information. Recent studies have explored various mechanisms for achieving quantum nonreciprocity, including dynamic modulation and parametric interactions in superconducting devices  (?, ?). Observations of superconducting nonreciprocity through diamagnetic effects and engineered circuit asymmetries offer robust platforms for nonreciprocal quantum devices (?, ?, ?, ?). Traditional directional absorbers and electronic components, such as varactors, transistors, and diodes, face limitations and introduce noise at millikelvin temperatures typical of superconducting quantum technologies. To overcome these challenges, we propose a spatiotemporally modulated superconducting metasurface that exhibits photon-blockade-analogue nonreciprocal absorption. By leveraging the unique properties of superconducting materials and dynamically modulating the metasurface structure, we achieve directional absorption where incident waves are selectively absorbed or transmitted based on their direction of incidence. This classical analogue of photon blockade, where forward-traveling waves undergo resonant harmonic conversion and absorption while backward-traveling waves transmit freely, offers new capabilities for controlling electromagnetic wave propagation in quantum-compatible platforms. The approach presents significant opportunities for advancing nonreciprocal devices and photon management in millikelvin-temperature quantum systems.
Our approach involves nonlinear space-time-modulated metasurfaces incorporating gate-controlled Josephson field-effect transistors (JoFETs)  (?, ?, ?). These metasurfaces deliver highly efficient directional absorption while maintaining superior performance in superconducting quantum applications  (?). By enabling the miniaturization and integration of directional absorption components, our method preserves quantum state integrity and represents a significant leap forward in superconducting quantum technologies. The metasurface, featuring cascaded space-time-varying JoFETs, exploits the quantum mechanical phenomenon of supercurrent flow without resistance  (?, ?). This design harnesses the nonlinearity of JoFETs and their spatial and temporal modulation capabilities to advance directional absorption and wave reflection. Our approach achieves efficient directional absorption with simultaneous amplification, a rare combination in conventional devices. It addresses inefficiencies and spurious signal generation, paving the way for high-performance signal processing applications.
@@
Furthermore, our metasurface demonstrates nonreciprocal behavior, with different absorption responses for incident waves from opposing directions, emphasizing the unidirectional properties imparted by the spatiotemporal modulation (?, ?, ?, ?) and time-periodic devices (?, ?, ?, ?, ?, ?, ?, ?). Utilizing analytical techniques such as Bloch-Floquet solutions, we provide a comprehensive understanding of the wave scattering phenomena from our metasurface, validating its design and offering insights for further optimization. Unlike linear systems constrained by phase-matching and dispersion, our nonlinear superconducting spatiotemporal media achieve superior performance in a compact form, making them ideal for millikelvin-temperature superconducting quantum technologies.
1 Photon-Blockade Analogue Mechanism for Nonreciprocal Absorption
@@
Figure 1: Nonreciprocal absorption in superconducting spatiotemporal metasurfaces, where the space-time modulation and incident wave share the same frequency ($\omega_{\text{s}}=\omega_{0}$). The system comprises space-time-modulated nonlinear unit cells coupled together. a) Incidence from the left port ($L_{\text{in}}$): Energy transition to higher states occur due to the strong interaction and one-way coherency with the nonlinear left-to-right traveling space-time modulation, resulting in zero transmission to the right port ($R_{\text{out}}=0$). b) Illustration of the physical mechanism for incidence from the right port ($R_{\text{in}}$): Free passage of wave to the left port ($L_{\text{out}}=R_{\text{in}}$) occurs due to the lack of transition to higher energy states and weak interaction with the opposite-direction traveling nonlinear space-time modulation, resulting in full transmission of waves.
Consider waves with frequency $\omega_{0}$ incident on a superconducting slab with a space-time periodic current density $J(z,t)$, modulated at a temporal frequency $\omega_{\text{s}}$, where $\omega_{\text{s}}=\omega_{0}$. We demonstrate that the temporal coherence between the incident wave and the periodic modulation of the superconducting slab induces a nonreciprocal transmission. This phenomenon arises because the energy associated with the space-time modulation effectively inhibits the transmission of subsequent photons. Instead, the incident photons transition to higher energy states within the slab and are absorbed, preventing their exit. This results in a one-way quantum absorption process, facilitated by the coherent spatiotemporal modulation of the superconducting slab.